OpenStudy (anonymous):

Evaluate the indefinite integral of: 1 ---- 1 + (cosx)^2 I would like only a hint(s) please.

OpenStudy (anonymous):

U-sub, integ by parts.... ?

OpenStudy (anonymous):

Any ideas Zarkon?

OpenStudy (anonymous):

Use a trig identity to make it possible

OpenStudy (anonymous):

You mean in the numerator? I would get sin^2(x) + cos^2(x)

OpenStudy (anonymous):

But how would that help at all?

OpenStudy (anonymous):

There is another identity, should it be revealed to you?

OpenStudy (anonymous):

Wow who knew a question could be so popular

OpenStudy (ash2326):

\[\int \frac{1}{1+\cos^2 x} dx\] multiply numerator and denominator by \(\sec^2 x\) \[\int \frac{\sec^2 x}{1+\sec^2x} dx\] \[\int \frac{\sec^2 x}{\tan^2x } dx\] put tan x= t \[\int \frac{dt}{t^2} \]

OpenStudy (anonymous):

Wow Ash, you got some ingenuity there!

OpenStudy (anonymous):

Is there any other way?

OpenStudy (zarkon):

\[\frac{1}{1+\cos^2(x)}=\frac{\sec^2(x)}{1+\sec^2(x)}=...\] stole my thunder ;)

OpenStudy (zarkon):

though...you hace some mistakes

OpenStudy (anonymous):

Wow. I had no idea. Ok I will remember that: multiple by the inverse identity. Check.

OpenStudy (ash2326):

@zarkon Where did I make mistakes?:D

OpenStudy (zarkon):

\[1+\sec^2(x)=1+1+\tan^2(x)=2+\tan^2(x)=\frac{1}{2}\left(1+\frac{\tan^2(x)}{2}\right)\]

OpenStudy (ash2326):

@Zarkon Oh thanks for pointing out \[\int \frac{ \sec^2 x}{2+\tan^2 x} dx\]

OpenStudy (zarkon):

\[1+\sec^2(x)=1+1+\tan^2(x)=2+\tan^2(x)=2\left(1+\frac{\tan^2(x)}{2}\right)\]

OpenStudy (zarkon):

had a little typo myself ;)

OpenStudy (anonymous):

Where did you get that 2 from?

OpenStudy (ash2326):

@QRAwarrior Sorry made a mistake:( Did you get it? \[\sec^2 x= 1+ \tan^2 x\] \[1+\sec^2 x=1+ 1+ \tan^2 x\]

OpenStudy (anonymous):

Got it.

OpenStudy (anonymous):

Wow are you people mathematicians or what? That was brilliant.

OpenStudy (ash2326):

@Zarkon is a real mathematician.:D Glad that we have him here:D